OpenStudy (laddiusmaximus):

find the derivative of the function using the definition of the derivative g(x)=sqrt(9-x) so far I have lim x-->0 =f(sqrt(9-(x-h))-f(sqrt(9-x)/h

zepdrix (zepdrix):

\[\large g(x)=\sqrt{9-x}\]\[\large g(x+h)=\sqrt{9-x-h}\]

zepdrix (zepdrix):

\[\large \lim_{h \rightarrow 0}\frac{ g(x+h)-g(x) }{ h }=\lim_{h \rightarrow 0}\frac{ \sqrt{9-x-h}-\sqrt{9-x} }{ h }\]